Calculate the value of the translational partition function of O2 at 1000 K for a volume of 1 m3. Submit your answer multiplied by 10−32 (i.e. if you calculate 4.2e33, input 42).
1
Expert's answer
2020-06-12T07:28:50-0400

"q_T=\frac{V}{\lambda^3}=\frac{V}{(\frac{h}{(2\pi mk_bT)^\frac{1}{2}})^3}"

"=\frac{1}{(\frac{6.626\times10^{-34}}{(2\times3.14\times0.032\times1.38\times10^{-23}\times 1000)^{0.5}})^{3}}"

"=1.693\times10^{34}"

Hence answer will be "169"